OpenStudy (anonymous):

solve. (x+4)(x-18)(x+2 >0

hero (hero):

Saifoo, help him

OpenStudy (anonymous):

(x + 4 )(x - 18) (x +2)> 0 Draw a Number Line <------------------------------------------|------------------------|--------------------------------|--------------------> \(-\infty\) -4 -2 18 \(\infty\)

OpenStudy (anonymous):

Now, check the intervals in which the function is greater than zero \(- - - - -\) \( +++ \) \(- - - - \) \(+++\) <------------------------------------------|------------------------|--------------------------------|--------------------> \(-\infty\) -4 -2 18 \(\infty\)

OpenStudy (anonymous):

which make the solution \[x \in (-4,-2)\cup(18,\infty)\]